1mole of NH3 gas(gamma=1.33) at 27°C is allowed to expand adiabatically reversibly so that the final volume becomes 8 times. What will be the work done?

Grade:11

2 Answers

Chhavi Jain
6 years ago
Dear student,
 
Given : Rev. adiabatic expansion(q=0) , Vf =8 V  , gamma= 1.33 , n = 1 , T= 300K 
              
                        using TVγ-1 = constant
                              
                                  T/ T= (V2 / V1 ) ^ (γ-1)
                                   300/ T2 =  8 ^ ( 1.33-1) 
 
                                    T= 150 K ( approx )
 
Now, using γ = Cp / Cv
              
                  1.33 = 4R / 3R = Cp /C
So, by compairing both sides we can say that :    Cv = 3R
 
Using : ΔE= q +w
          As,   q=0    ;  ΔE= w
 
                               w = nCvΔT
                                w= 1 (3R) ( T-T1 )
                                w= 3R ( 150-300 )
                                w= -150R 
 
 
Chemistry Consultant
Askiitians 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Given : Rev. adiabatic expansion(q=0) , Vf=8 Vi , gamma= 1.33 , n = 1 , T= 300K
using TVγ-1= constant
T1/ T2= (V2 / V1)^(γ-1)
300/ T2= 8^ ( 1.33-1)
T2= 150 K ( approx )
Now, usingγ = Cp/ Cv
1.33 = 4R / 3R = Cp/Cv
So, by compairing both sides we can say that : Cv= 3R
Using : ΔE= q +w
As, q=0 ;ΔE= w
w = nCvΔT
w= 1 (3R) ( T2-T1)
w= 3R ( 150-300 )
w= -150R

Thanks and Regards